OpenStudy (butterflyhope):

A shopper bought four shirts, all at the same price, and spent $75 on ties. The total cost of $155. What was the cost per shirt? I need alot of help today, wow xD Thanks!

OpenStudy (anonymous):

let x be the price of 1 shirt so he bought 4 shirts therefore 4x also he bought ties cost 75 till now the cost of shirts and ties is 4x+75 the total cost is mentioned 155$ so 4x+75=155 get x ... 4x=155-75 4x=80 x= 80/4 x=20

OpenStudy (anonymous):

hope this is useful

OpenStudy (butterflyhope):

Thank you :)

OpenStudy (anonymous):

welcome :)